Which stock exchange is a “virtual exchange”?

      I. London stock exchange

             II. New York Stock exchange

III. Tokyo stock exchange

            IV. Over-the-counter market

I and II only

III and IV only

I only

  

   IV only

Kensington Company stock was selling at $132 a share when Charlotte sold 300 shares of the stock short. Today Charlotte bought 300 shares of the same stock at a price of $140 per share to cover her position. Ignoring trading costs, what is the dollar return on Charlotte's investment?

$2,400

-$800

$800

$-2,400

Tom is actively adding more diversification to his portfolio. Because of this diversification increase, the firm-specific risk of his portfolio will move toward:

0.

1.

infinity.

(n – 1) × n.

The widely used measure for measuring portfolio performance devised by Treynor focuses on:

non-diversifiable risk.

diversifiable risk.

the standard deviation of the portfolio.

total risk.

Mila, a muni bond portfolio manager, would inform you that which of the following are true regarding municipal bonds?

I) A municipal bond is a debt obligation issued by state or local governments.

II) A municipal bond is a debt obligation issued by the federal government.

III) The interest income from a municipal bond is exempt from federal income taxation.

IV) The interest income from a municipal bond is exempt from state and local taxation in the issuing state.

I and II only

I and III only

I, II, and III only

I, III, and IV only

I and IV only

If the present value of an investment's benefits equals the present value of the investment's costs, then the investor would earn a:

return equal to the discount rate.

negative rate of return.

0% rate of return.

return greater than the discount rate.

Trent's portfolio has a beta of 1.37 and earns a return of 14%. Sophie's portfolio has a beta of 0.8 and earns a return of 11%. The risk-free rate is 3% and the expected rate of return on the market is 12%. According to the Jensen's measure:

Trent has the better portfolio.

Sophie has the better portfolio.

the portfolios are equally desirable.

the answer depends on Trent and Sophie's risk tolerance.

You are taking the CFA Level II exam and encounter the following question: If the risk-free rate is 0.06 and the expected market rate of return is 0.12, then according to the capital asset pricing model (CAPM), the expected rate of return on security X with a beta of 1.2 is equal to:

0.06

0.144

0.12

0.132

0.18
Answer #1

1.
I and II only

2.
=300*(140-132)
=2400

3.
0

4.
non diversifiable risk

5.
I, III, and IV only

6.
return equal to the discount rate.

7.
Sophie has the better portfolio.
Jensen's measure=expected return-risk free rate-beta*(market return-risk free rate)
Trent=14%-3%-1.37*(12%-3%)=-1.33%
Sophie=11%-3%-0.8*(12%-3%)=0.80%

8.
=risk free rate+beta*(market return-risk free rate)
=0.06+1.2*(0.12-0.06)
=0.132